3 × 10000 + 7 × 1000 + 9 × 100 + 0 ×10 + 4 is the same as

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3 × 10000 = 30000
The place value is 30000
7 × 1000 = 7000
The place value is 7000
9 × 100 = 900
The place value is 900
0 × 10 = 0
4= 4
Thus, the sum of all place value is = 30000 + 7000 + 900 + 0 + 4

= 37904

Thanks